A charged particle in moving under the influence of a point nucleus. Find the orbit of the particle and periodic time in the case of an elliptical orbit.


1
Expert's answer
2021-01-23T10:18:48-0500

Answer

A charged particle(let electron) in moving under the influence of a point nucleus.

Then central force is given

"F=\frac{-ze^2}{r^2}" ";k=ze^2"

E="\frac{-k}{2a}=\frac{-ze^2}{2a}"

Eccentricity

"\epsilon=\sqrt{1+\frac{2Eh^2}{mz^2e^4}}=\sqrt{1+\frac{h^2}{amz^2e^2}}"

This shows elliptical path.

Now time period

"T=\frac{4\pi ^2a^2m}{k}=\frac{2m}{e}\sqrt{\frac{ma^2}{z}}"